OpenStudy (anonymous):

Find g(1) if g(x) = x2 + 1. 2 3 4

OpenStudy (anonymous):

g(1) means x =1 so substitute 1 for x and solve g(1) = 1^2 +1

OpenStudy (anonymous):

i don't understand what that means..

OpenStudy (anonymous):

g(X)=x2+1. now put 1 in place of x in the equation. g(1)=2*1+1=2+1=3

OpenStudy (anonymous):

ohhhh, thank you!!!!!!!

OpenStudy (anonymous):

g(1) = 2 + 1. g(1) = 3

OpenStudy (anonymous):

u r welcum.....